what are four consecutive integers such that the 3rd number added to double the fourth number is four times the second number?
Answer by checkley77(12844)

Let x, x+1, x+2, & x+3 be the 4 integers.
(x+2)+2(x+3)=4(x+1)
x+2+2x+6=4x+4
3x+8=4x+4
3x-4x=4-8
-x=-4
x=4 the smallest integer.
4+1=5 the second integer.
4+2=6 the third integer.
4+3-7 the largest integer.
Proof:
6+2*7=4*5
6+14=20
20=20